OpenStudy (anonymous):

17 fifthsqrt3x^2-15 fifthroot 3x^2

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

Is this the question? \(\LARGE 17 \sqrt[5]{3x^2} -15\sqrt[5]{3x^2} \)

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Notice that the two roots are the same. They are both the fifth root of 3x^2. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i combine something.

OpenStudy (anonymous):

That means you have like terms. You can combine like terms. What is 17a - 15a = ?

OpenStudy (solomonzelman):

Don't say `fifth square root of...` it is either just `fifth root` or `square root` .

OpenStudy (anonymous):

oh sorry.. @SolomonZelman

OpenStudy (anonymous):

2 @MaxwellSmart

OpenStudy (solomonzelman):

It's okay, I am just saying, because when you say this you scare off some very smart people like tkhunny. he, for example, if he sees that he would be very critical about it... but nothing personal or bad... just explaining .

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

17a - 15a = 2a 17 oranges - 15 oranges = 2 oranges 17 math students - 15 math students = 2 math students In the same way \(\LARGE 17 \sqrt[5]{3x^2} -15\sqrt[5]{3x^2} \) \(\LARGE =2\sqrt[5]{3x^2}\)

OpenStudy (anonymous):

haha okay! great! @MaxwellSmart

OpenStudy (anonymous):

i have one more if any of you care to do it... @MaxwellSmart @SolomonZelman

OpenStudy (anonymous):

Yes, but pls start a new post. One question per post.

OpenStudy (anonymous):

okay